### v2

# A parameterized linear formulation of the integer hull

Abstract
Let $A\in\mathbb{Z}^{m\times n}$ be an integer matrix with components bounded by $\Delta$ in absolute value. Cook et al. (1986) have shown that there exists a universal matrix $B\in\mathbb{Z}^{m^{\prime}\times n}$ with the following property: For each $b\in\mathbb{Z}^{m}$, there exists $t\in\mathbb{Z}^{m^{\prime}}$ such that the integer hull of the polyhedron $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ is described by $P_{I}=\{x\in\mathbb{R}^{n}\colon Bx\leq t\}$. Our main result is that $t$ is an affine function of $b$ as long as $b$ is from a fixed equivalence class of the lattice $D\cdot\mathbb{Z}^{m}$. Here $D\in\mathbb{N}$ is a number that depends on $n$ and $\Delta$ only. Furthermore, $D$ as well as the matrix $B$ can be computed in time depending on $\Delta$ and $n$ only. An application of this result is the solution of an open problem posed by Cslovjecsek et al. (SODA 2024) concerning the complexity of 2-stage-stochastic integer programming problems. The main tool of our proof is the classical theory of Gomory-Chvátal cutting planes and the elementary closure of rational polyhedra.
1 Introduction
An integer program is an optimization problem of the form
$\max\{c^{T}x\colon Ax\leq b,\,x\in\mathbb{Z}^{n}\},$ (1)
where $c\in\mathbb{Z}^{n}$, $A\in\mathbb{Z}^{m\times n}$ and $b\in\mathbb{Z}^{m}$. Many optimization problems can be modeled and solved as an integer program, see, e.g. [18, 22, 6, 21]. Unlike linear programming, which can be solved in polynomial time [16], integer programming is known to be NP-complete [3].
An important special case arises if the polyhedron $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ is integral, i.e., if $P$ is equal to its integer hull $P_{I}=\operatorname{conv}(\mathbb{Z}^{n}\cap P)$. Here $\operatorname{conv}(X)$ denotes the convex hull of a set $X\subseteq\mathbb{R}^{n}$. In this case, the integer program (1) can be solved in polynomial time with linear programming and integer linear algebra [21]. For example, if the matrix $A$ is totally unimodular, i.e., if each sub-determinant of $A$ is 00 or ±$\pm 1$, then $P$ is integral for each right-hand-side vector $b\in\mathbb{Z}^{m}$. Here, the notion sub-determinant of $A$ refers to the determinant of some square sub-matrix of $A$. Efficient algorithms for integer programs defined by matrices $A$ with a fixed sub-determinant bound is an active area of research [2, 11, 1].
In the following we consider polyhedra defined by the matrix $A\in\mathbb{Z}^{m\times n}$ with varying right-hand-side $b\in\mathbb{Z}^{m}$ and denote the thereby parameterized polyhedron by $P(b)=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$. Cook, Gerards, Schrijver and Tardos [8] have shown that there exists a universal integral matrix $M\in\mathbb{Z}^{m^{\prime}\times n}$ that depends only on the matrix $A$ with the following property:
For each right-hand-side $b\in\mathbb{Z}^{m}$ there exists a right-hand-side $t\in\mathbb{Z}^{m^{\prime}}$ with
$P(b)_{I}=\{x\in\mathbb{R}^{n}\colon Mx\leq t\}.$ (2)
Furthermore, Cook et al. show that $\|M\|_{\infty}\leq n^{2n}\delta^{n}$. Here $\delta$ is an upper bound on the sub-determinants of $A$ and $\|M\|_{\infty}$ denotes the largest absolute value of a component of $M$.
Contribution
We prove that the right-hand-side $t\in\mathbb{Z}^{m^{\prime}}$ in (2) depends linearly on the right hand side $b\in\mathbb{Z}^{m}$ for all $b$ belonging to the same equivalence class of a certain lattice that depends on $\Delta$ and $n$ only. We make the natural assumption that $A\in\mathbb{Z}^{m\times n}$ does not have repeating rows. Recall that the number $m$ of rows is bounded by $m\leq(2\Delta+1)^{n}$ in this case. Our main contribution is the following.
Theorem 1.
Let $A\in\mathbb{Z}^{m\times n}$ with non-repeating rows and $\|A\|_{\infty}\leq\Delta$. There exists a $D\in\mathbb{N}$ depending on $n$ and $\Delta$ such that, for each $r\in\{0,\dots,D-1\}^{m}$, there exist $B_{r}\in\mathbb{Z}^{m^{\prime}\times n}$, $C_{r}\in\mathbb{Z}^{m^{\prime}\times m}$ and $f_{r}\in\mathbb{Z}^{m^{\prime}}$ such that the following holds:
For each $b\in\mathbb{Z}^{m}$ with $b-r\in{D\cdot\mathbb{Z}^{m}}$, one has
$P(b)_{I}=\left\{x\in\mathbb{R}^{n}\colon B_{r}x\leq f_{r}+C_{r}b\right\}.$ (3)
Furthermore, the number $D\in\mathbb{N}$, the matrices $B_{r}$ and $C_{r}$, as well as the vector $f_{r}$ can be computed in time depending on $\Delta$ and $n$ only.
Remark 1.
The condition that $A\in\mathbb{Z}^{m\times n}$ has non-repeating rows can also be dropped. In this case, $D$, $B_{r}$ and $C_{r}$, as well as the vector $f_{r}$, as well as the running time depend on $\Delta$, $n$, and $m$.
Applications
We describe three almost immediate applications of Theorem 1.
i) A 2-stage stochastic integer programming problem is an integer program (1) where the constraints can be described as maximizing $c^{T}x+\sum_{i=1}^{n}d_{i}^{T}y_{i}$ subject to
$U_{i}x+V_{i}y_{i}=b_{i},\,i=1,\dots,n\,\text{ and }\,x,y_{1},\dots,y_{n}\in\mathbb{Z}_{\geq 0}^{k}.$
If $\Delta$ denotes the largest absolute value of an entry of the integral constraint matrices $U_{i},V_{i}\in\mathbb{Z}^{k\times k}$, then the problem can be solved in time $f(\Delta,k)$-times a polynomial in the input encoding [14, 17, 9]. See also [15] for a doubly exponential lower bound on $f$. Recently Clovjecsek et al. [10] have proved that the integer feasibility problem of 2-stage stochastic integer programming can be solved in time $f(\Delta,k)$-times a polynomial in the input encoding. This time, $\Delta$ is a bound on the largest absolute value of the matrices $V_{i}$ only. The authors pose the open problem whether such a complexity bound follows for the full problem (optimization version) as well. We answer this question in the affirmative in Section 4.2.
ii) Clovjecsek et al. [10] provide an important structural result on the convexity of integer cones. In a nutshell, their result shows that the elements of an integer cone generated by a matrix $W\in\mathbb{Z}^{m\times n}$ are partitioned by a finite number of shifts of a lattice $\Lambda=D\cdot\mathbb{Z}^{m}$. More precisely, for each $r\in\{0,\dots,D-1\}^{m}$ there exists a polyhedron $Q_{r}\subseteq\mathbb{R}^{m}$ such that
$(r+\Lambda)\cap\operatorname{intcone}(W)=(r+\Lambda)\cap Q_{r}.$
Their original proof is fairly involved. It is implied by our main theorem very quickly as we lay out in Section 4.1.
iii) Finally, we turn our attention to 4-block integer programming problems for which there only exist polynomial-time algorithms, where the degree of the polynomial depends on the parameters $\Delta$ and $k$ [13]. We prove that in time $f(\Delta,k)$ times a polynomial of fixed degree one can at least find an almost feasible solution that violates the $O(1)$ many linking constraints by an additive constant depending on $k$ and $\Delta$. The details are in Section 4.3.
Proof idea
The proof of Theorem 1 is based on the classical theory of Gomory-Chvátal cutting planes and the elementary closure of polyhedra. We will review this theory in Section 2 at a level of detail that is necessary for us here. The reader shall also be referred to the seminal textbooks [21, 6].
However, the main idea can be easily described by resorting to the following main principles of the theory of cutting planes: The elementary closure of a rational polyhedron $P$ is again a rational polyhedron $P^{\prime}$ that satisfies $P_{I}\subseteq P^{\prime}\subseteq P$. The computation of $P^{\prime}$ is fixed-parameter-tractable in $\Delta$ and $n$. Furthermore, if $P^{(i)}$ denotes the result of $i$ successive applications of the closure operator, then the Chvátal rank of $P$ is the smallest $i\in\mathbb{N}$ with $P^{(i)}=P_{I}$. It is always finite and, for us very important, it is bounded by a function on $n$ and $\Delta$ [8]. Together, this indicates that, in order to prove our main result, it is enough to show some analogous statement for $P^{(i)}$ instead for $P_{I}$ directly. This is done in Theorem 5 and constitutes the heart of the proof.
2 Gomory-Chvátal cutting planes
Let $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ be a polyhedron where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$. For any $c\in\mathbb{Z}^{n}$ and $\delta\geq\max\{c^{T}x\colon Ax\leq b\}$, the inequality $c^{T}x\leq\lfloor\delta\rfloor$ is called a Gomory-Chvátal cutting plane [12, 5], or briefly cutting plane of $P$. It is valid for all integer points $x\in P\cap\mathbb{Z}^{n}$ and therefore also for the integer hull $P_{I}$ of $P$.
[…]
Figure 1: The valid inequality $-x_{1}+x_{2}\leq\delta$ yields the cutting plane $-x_{1}+x_{2}\leq\lfloor\delta\rfloor$.
For the ease of notation we will also write $c^{T}x\leq\delta$ for the set of points $\{x\in\mathbb{R}^{n}:c^{T}x\leq\delta\}$. The intersection of all cutting planes of $P$ is denoted by
$P^{\prime}=\bigcap_{\begin{subarray}{c}(c^{T}x\leq\delta)\supseteq P\\ c\in\mathbb{Z}^{n}\end{subarray}}\big{(}c^{T}x\leq\lfloor\delta\rfloor\big{)}.$ (4)
A cutting plane $c^{T}x\leq\lfloor\delta\rfloor$ follows from a simple inference rule that is based on linear programming duality. Duality implies that there exists a $\lambda\in\mathbb{R}^{m}_{\geq 0}$ such that
  1. i)
    $\lambda^{T}A=c^{T}$ and
  2. ii)
    $\lambda^{T}b\leq\delta$ hold.
Therefore, the elementary close can be described as follows
$P^{\prime}=\bigcap_{\begin{subarray}{c}\lambda\in\mathbb{R}^{m}_{\geq 0}\\ \lambda^{T}A\in\mathbb{Z}^{n}\end{subarray}}\big{(}(\lambda^{T}A)x\leq\lfloor\lambda^{T}b\rfloor\big{)}.$ (5)
If the polyhedron $P$ is rational, then the matrix $A$ in the representation $Ax\leq b$, as well as $b$ can be chosen to be integral, i.e., $A\in\mathbb{Z}^{m\times n}$, $b\in\mathbb{Z}^{m}$. If this is the case, then $P^{\prime}\subseteq P$ and $P^{\prime}$ is a rational polyhedron as well [20]. We repeat the argument below and provide a useful bound on the $\ell_{\infty}$-norm of the left-hand-side vector $c$ of a non-redundant cutting plane $c^{T}x\leq\lfloor\delta$$\rfloor$.
Theorem 2 (Schrijver [20]).
Let $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ be a polyhedron with $A\in\mathbb{Z}^{m\times n}$, $\|A\|_{\infty}\leq\Delta$ and $b\in\mathbb{Z}^{m}$. Then
$P^{\prime}=\bigcap_{\begin{subarray}{c}\lambda\in[0,1)^{m}\\ \lambda^{T}A\in\mathbb{Z}^{n}\end{subarray}}\big{(}(\lambda^{T}A)x\leq\lfloor\lambda^{T}b\rfloor\big{)}.$
In particular, in (4), $P^{\prime}$ is described by all cutting planes $c^{T}x\leq\lfloor\delta\rfloor$ with $c\in\mathbb{Z}^{n}$ and $\|c\|_{\infty}\leq n\cdot\Delta$.
Proof.
Consider a cutting plane $c^{T}x\leq\lfloor\delta\rfloor$ derived as $(\lambda^{T}A)x\leq\lfloor\lambda^{T}b\rfloor$ with $\lambda\in\mathbb{R}_{\geq 0}^{m}$, $\lambda^{T}A=c^{T}$ and $\lambda^{T}b\leq\delta$.
The inequality $(\lambda-\lfloor\lambda\rfloor)^{T}Ax\leq\lfloor(\lambda-\lfloor\lambda\rfloor)^{T}b\rfloor$ is again a cutting plane, since $(\lambda-\lfloor\lambda\rfloor)^{T}A\in\mathbb{Z}^{n}$. We have
$\lambda^{T}Ax=(\lambda-\lfloor\lambda\rfloor)^{T}Ax+\lfloor\lambda\rfloor^{T}Ax\leq\lfloor(\lambda-\lfloor\lambda\rfloor)^{T}b\rfloor+\lfloor\lambda\rfloor^{T}b=\lfloor\lambda^{T}b\rfloor.$
This implies that the inequality $(\lambda^{T}A)x\leq\lfloor\lambda^{T}b\rfloor$ is linearly implied by the system $Ax\leq b$ and the inequality $(\lambda-\lfloor\lambda\rfloor)^{T}Ax\leq\lfloor(\lambda-\lfloor\lambda\rfloor)^{T}b\rfloor$.
Furthermore, we can assume that $\lambda$ is an optimal vertex solution to the linear program
$\min\{\lambda^{T}b\colon\lambda^{T}A=c^{T},\,\lambda\in\mathbb{R}^{m}_{\geq 0}\},$ (6)
which implies that $\lambda$ has only $n$ nonzero components. From there, it follows that
$\|(\lambda-\lfloor\lambda\rfloor)^{T}A\|_{\infty}\leq n\cdot\Delta.$
By applying the closure operator $i$-times successively, one obtains $P^{(i)}$, the $i$-th elementary closure of $P\subseteq\mathbb{R}^{n}$. The next corollary is immediate.
Corollary 3.
Let $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ be a rational polyhedron with $A\in\mathbb{Z}^{m\times n}$, $\|A\|_{\infty}\leq\Delta$ and $b\in\mathbb{Z}^{m}$. Let $i\in\mathbb{Z}_{\geq 0}$. Then $P^{(i)}$ is described by a system of inequalities
$P^{(i)}=\{x\in\mathbb{R}^{n}\colon Cx\leq d\},$
$C\in\mathbb{Z}^{m^{\prime}\times n}$, $d\in\mathbb{Z}^{m^{\prime}}$ with $\|C\|_{\infty}\leq n^{i}\Delta$.
The rank of rational polyhedra
The Chvátal rank [20] of a polyhedron $P$ is the smallest natural number $i\in\mathbb{Z}_{\geq 0}$ such that $P^{(i)}=P_{I}$. Schrijver [20] has shown that the Chvátal rank of a rational polyhedron is always finite.
Cook et al. [7] showed that the Chvátal rank of integer empty polyhedra is bounded by a function on the dimension. More precisely, they show that, the rank of $P\subseteq\mathbb{R}^{n}$ with $P\cap\mathbb{Z}^{n}=\emptyset$ is bounded by a function $g(n)$ that satisfies the recursion
$g(n)\leq\omega(n)(1+g(n-1))+1,$ (7)
where $\omega(n)$ is the flatness constant in dimension $n$. The best known bound on $\omega(n)$ is $O(n\log^{3}(2n))$, see [19], which gives a bound of $g(n)\leq n^{O(n)}$.
Most relevant for us is another result of Cook et al. [8] that states that the rank can be bounded by $\Delta$ and $n$. It follows from a proximity theorem, the bound on $g(n)$ (7) and the bound on the $\ell_{\infty}$-norm of the facet normal-vectors of $P_{I}$, see also [21, Theorem 23.4]. More precisely, they show that the rank of $P(b)$ is bounded by
$\max\left\{g(n),n^{2n+2}\delta^{n+1}+1+n^{2n+2}\delta^{n+1}g(n-1)\right\}$ (8)
where $g(n)$ is as in (7) and $\delta$ is an upper bound on the largest sub-determinant of $A$. With the Hadamard bound on $\delta$ one can see that there exists a constant $C\in\mathbb{N}$ such that (8) is bounded from above by $(n\cdot\Delta)^{C\cdot n^{2}}$. We summarize this in the following theorem.
Theorem 4 (Cook, Gerards, Schrijver and Tardos [8]).
Let $P=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}$ be a rational polyhedron with $A\in\mathbb{Z}^{m\times n}$, $\|A\|_{\infty}\leq\Delta$ and $b\in\mathbb{Z}^{m}$. There exists a function $\operatorname{Rank}(\Delta,n)$ such that
$P^{(i)}=P_{I}$
for each $i\geq\operatorname{Rank}(\Delta,n)$. Moreover, there exists a constant $C\in\mathbb{N}$ such that
$\operatorname{Rank}(\Delta,n)\leq(n\cdot\Delta)^{C\cdot n^{2}}.$ (9)
3 Proof of the main theorem
This section contains the proof of Theorem 1. Throughout, we assume that $A\in\mathbb{Z}^{m\times n}$ satisfies $\|A\|_{\infty}\leq\Delta$. The strategy is to show an analogous statement to Theorem 1 for the $i$-th elementary closure. Let us start with the first Gomory-Chvátal closure.
3.1 Linearity of the first elementary closure
Consider a cutting plane $c^{T}x\leq\lfloor\delta\rfloor$, where $c\in\mathbb{Z}^{n}$. If this cutting plane is non-redundant, then there exists an optimal vertex solution $\lambda\in\mathbb{R}^{m}_{\geq 0}$ of the linear program (6) with $\delta\geq\lambda^{T}b$. This means that there are at most $n$ linearly independent rows of $A$, indexed by $B\subseteq\{1,\dots,m\}$ such that $\lambda_{B}$ is the unique solution of the linear equation
$\lambda_{B}^{T}A_{B}=c^{T},$
with all other components of $\lambda$ equal to zero. By deleting linearly dependent columns of $A_{B}$ and the corresponding components of $c$, this shows that there is a $|B|\times|B|$ non-singular sub-matrix $\overline{A}$ of $A$ with
$\lambda_{B}^{T}\overline{A}\in\mathbb{Z}^{|B|}.$
Cramer’s rule implies then that $\lambda=\mu/D$ where $\mu\in\mathbb{Z}_{\geq 0}^{m}$, where $D\in\mathbb{N}$ is a multiple of all sub-determinants of $A$.
Furthermore, we have seen in Theorem 2 that we can replace $\lambda\in\mathbb{R}_{\geq 0}^{m}$ by $\lambda-\lfloor\lambda\rfloor$. Thus we have
$\lambda=\mu/D\,\text{ with }\,\mu\in\{0,\ldots,D-1\}^{m}.$
A Gomory-Chvátal cut $(\mu/D)^{T}Ax\leq\lfloor(\mu/D)^{T}b\rfloor$ with $\mu\in\{0,\ldots,D-1\}^{m}$ and $\mu^{T}A\equiv 0\pmod{D}$ is called $\mathrm{mod-}D$ cut [4]. In other words, if $D\in\mathbb{N}$ is an integer multiple of the sub-determinants of $A$, then the set of $\mathrm{mod-}D$ cuts contains all non-redundant cutting planes.
Let us now fix a $\mu\in\{0,\ldots,D-1\}^{m}$ with $\mu^{T}A\equiv 0\pmod{D}$ as well as a remainder $r\in\{0,\ldots,D-1\}^{m}$ and consider $b\in\mathbb{Z}^{m}$ with $b\equiv r\pmod{D}$. The cutting plane derived by $\lambda=\mu/D$ can be written as
$\begin{array}[]{rcl}\underbrace{\left({\mu}/{D}\right)^{T}A}_{\text{row of }B_{r}}x&\leq&\lfloor(\mu/D)^{T}b\rfloor\\ &=&\underbrace{\lfloor(\mu^{T}r)/D\rfloor-(\mu^{T}r)/D}_{\text{component of }f_{r}}+\underbrace{(\mu^{T}/D)}_{\text{row of }C_{r}}b,\end{array}$ (10)
where the equation above follows from the fact that $\mu^{T}(b-r)/D$ is an integer. We conclude that the right-hand-side of the cut induced by $\lambda=\mu/D$ is indeed is linear in $b$.
For fixed $r\in\{0,\ldots,D-1\}^{m}$ we therefore enumerate all $\mu\in\{0,\ldots,D-1\}^{m}$ with $\mu^{T}A\equiv 0\pmod{D}$, fill the corresponding rows/components of $B_{r},C_{r}$ and $f_{r}$ as in (10) and append the original inequalities $Ax\leq b$.
Remark 2.
The above constructed $C_{r}$ and $f_{r}$ are rational and not-necessarily integral. In order to be conform with the notation in Theorem 1 we thus scale with the least common multiple of the denominators to obtain integral data.
3.2 Linearity of the $i$-th elementary closure
We now proceed with the general statement and its proof. Recall that the Chvátal rank of $P(b)$ is bounded by $\operatorname{Rank}(\Delta,n)$ and that each $P(b)^{(i)}$ can be described by a constraint matrix with infinity norm bounded by $n^{\max\{i,\operatorname{Rank}(\Delta,n)\}}\Delta$. We define $D\in\mathbb{N}$ to be the least common multiple of nonzero sub-determinants of $n\times n$-integer matrices with entries bounded by $n^{\operatorname{Rank}(\Delta,n)}\cdot\Delta$ in absolute value. In other words, let
$D=\operatorname{lcm}\left\{\det(C)\colon C\in\mathbb{Z}^{k\times k},\,k\leq n,\,\|C\|_{\infty}\leq n^{\operatorname{Rank}(\Delta,n)}\cdot\Delta,\,\det(C)\neq 0\right\}.$ (11)
It follows that $P^{(i+1)}$ is defined by $\mathrm{mod-}D$-cuts derived from the description of $P^{(i)}$ as in Corollary 3. In other words, $D$ is the largest modulus that is ever necessary in the derivation of a valid cutting plane. Theorem 1 is now an immediately corollary of the following statement concerning $P^{(i)}$.
Theorem 5.
Let $A\in\mathbb{Z}^{m\times n}$ with non-repeating rows and $\|A\|_{\infty}\leq\Delta$ and let $D$ be as in (11). For each $i\in\mathbb{Z}_{\geq 0}$ and $r\in\{0,\dots,D^{i}-1\}^{n}$, there exist $B_{r}\in\mathbb{Z}^{m^{\prime}\times n}$, $C_{r}\in\mathbb{Z}^{m^{\prime}\times m}$ and $f_{r}\in\mathbb{Z}^{m^{\prime}}$ such that for each $b\in\mathbb{Z}^{m}$ with $b-r\in D^{i}\cdot\mathbb{Z}^{m}$, one has
$P(b)^{(i)}=\left\{x\in\mathbb{R}^{n}\colon B_{r}\,x\leq f_{r}+\frac{C_{r}(b-r)}{D^{i}}\right\}.$ (12)
Furthermore, $B_{r}$ satisfies $\|B_{r}\|_{\infty}\leq n^{i}\Delta$.
Proof.
We prove the assertion by induction on $i$. For $i=0$, one has
$P(b)^{(0)}=P(b)=\{x\in\mathbb{R}^{n}\colon Ax\leq b\}.$
Each $b\in\mathbb{Z}^{m}$ is congruent to $\bm{0}$ modulo ${1=D^{0}}$. We set, for $r=\bm{0}$,
$B_{r}=A,\,f_{r}=\bm{0},\text{ and }C_{r}=I.$
We assume that the assertion is true for $i\in\mathbb{Z}_{\geq 0}$ and show that it is true for $i+1$ as well. If $i\geq\operatorname{Rank}(\Delta,n)$, then $P(b)^{(i+1)}=P(b)^{(i)}$ and we are done.
Otherwise let $r\in\{0,\dots,D^{i+1}-1\}^{m}$ and let $b\equiv r\pmod{D^{(i+1)}}$. By the induction hypothesis,
$P(b)^{(i)}=\left\{x\in\mathbb{R}^{n}\colon B_{r}x\leq f_{r}+\frac{C_{r}(b-r)}{D^{i}}\right\}$
for each $b$ with $b\equiv r\pmod{D^{i}}$ and in particular for each $b$ with $b\equiv r\pmod{D^{i+1}}$. The absolute values of the entries of $B_{r}\in\mathbb{Z}^{m^{\prime}\times n}$ are bounded from above by $n^{i}\Delta$.
We now construct the cutting planes of $P(b)^{(i)}$. The discussion above implies that a non-redundant cutting plane of $P(b)^{(i)}$ is of the form
$(\mu/D)^{T}B_{r}x\leq\left\lfloor(\mu/D)^{T}\left(f_{r}+\frac{C_{r}(b-r)}{D^{i}}\right)\right\rfloor,$ (13)
where $\mu\in\{0,\dots,D-1\}^{m^{\prime}}$ is of support at most $n$ and $\mu^{T}B_{r}\equiv 0\pmod{D}$. We wish to describe integer matrices $B_{r}^{\prime}$, $C_{r}^{\prime}$ and an integer vector $f_{r}^{\prime}$ such that
$P(b)^{(i+1)}=\left\{x\in\mathbb{R}^{n}\colon B_{r}^{\prime}x\leq f_{r}^{\prime}+\frac{C_{r}^{\prime}(b-r)}{D^{i+1}}\right\}$
Let $B_{r}^{\prime}$ be the integer matrix whose rows are comprised first, of the rows of $B_{r}$ and then of the left-hand-side vectors of cutting planes (13). Clearly, $\|B_{r}^{\prime}\|_{\infty}\leq n^{i+1}\Delta$. Since $b\equiv r\pmod{D^{i+1}}$, the right-hand-side of (13) is
$\left\lfloor(\mu/D)^{T}\left(f_{r}+\frac{C_{r}(b-r)}{D^{i}}\right)\right\rfloor$ $=$ $\left\lfloor(\mu/D)^{T}f_{r}+\frac{\mu^{T}C_{r}(b-r)}{D^{i+1}}\right\rfloor$
$=$ $\underbrace{\left\lfloor(\mu/D)^{T}f_{r}\right\rfloor}_{\text{ component of }f^{\prime}_{r}}+\frac{\overbrace{\mu^{T}C_{r}}^{\text{row of }C^{\prime}_{r}}(b-r)}{D^{i+1}},$
where the last equation follows from the fact that each component of $b-r$ is divisible by $D^{i+1}$ and therefore that
$\frac{\mu^{T}C_{r}(b-r)}{D^{i+1}}\in\mathbb{Z}.$
Remark 3.
  1. i)
    The proof of Theorem 5 contains a finite and constructive procedure to compute $D$, the matrices $B_{r},C_{r}$ and the vector $f_{r}$. The input to this procedure is the matrix $A\in\mathbb{Z}^{m\times n}$. Since $A$ does not have repeated rows, the running time of this procedure is a function in $\Delta$ and $n$ only.
  2. ii)
    The condition in Theorem 5 that $A\in\mathbb{Z}^{m\times n}$ has non-repeating rows can be dropped. In this case, we have a dependence on $n$, $m$ and $\Delta$.
  3. iii)
    Using the Hadamard bound and $\operatorname{Rank}(\Delta,n)\leq(n\cdot\Delta)^{C\cdot n^{2}}$ from Theorem 4, the number $D$ can be bounded from above by
    $D$$\leq$ $\operatorname{lcm}\left\{1,\dots,n^{{(n\cdot\Delta)}^{C^{\prime}\cdot n^{2}}}\right\}$
    $\leq$ $2^{n^{{n\cdot\Delta}^{C^{\prime\prime}\cdot n^{2}}}},$
    where $C^{\prime},C^{\prime\prime}\in\mathbb{N}_{+}$ are suitable constants. This bound is triple-exponential in $n$.
4 Applications
4.1 The convexity of integer cones
For a matrix $W\in\mathbb{Z}^{m\times n}$, we denote the integer cone as $\operatorname{intcone}(W)=\{Wx:x\in\mathbb{Z}_{\geq 0}^{n}\}$. An integer cone is a discrete set and in general it will have “holes”, i.e., integer points in the rational cone of $W$ that are not in the integer cone. But Cslovjecsek, Koutecký, Lassota, Pilipczuk and Polak [10] proved that integer cones are indeed convex in a discrete sense when intersected with shifts of certain sparse lattices.
Theorem 6 (Convexity of Integer Cones [10]).
Let $W\in\mathbb{Z}^{m\times n}$ with $\|W\|_{\infty}\leq\Delta$. Then there is a $D\in\mathbb{N}$ dependent only on $m$ and $\Delta$ so that for every $r\in\{0,\ldots,D-1\}^{m}$ there exists a polyhedron $Q_{r}\subseteq\mathbb{R}^{m}$ so that
$\Lambda_{r}\cap\operatorname{intcone}(W)=\Lambda_{r}\cap Q_{r}$
where $\Lambda_{r}=r+D\cdot\mathbb{Z}^{m}$.
The original proof of this result takes a substantial amount of work. We demonstrate that it is quickly implied by our main result.
Proof of Theorem 6.
We can assume that $W\in\mathbb{Z}^{m\times n}$ does not have repeated columns. Therefore, $n$ is bounded by $(2\cdot\Delta+1)^{m}$. Consider $P(b)=\{x\in\mathbb{R}^{n}:Wx=b,x\geq\bm{0}\}$. Observe that the condition $Wx=b$ and $x\geq\bm{0}$ can be formulated in the usual inequality standard-form as the conjunction of $Wx\leq b$, $-Wx\leq-b$ and $-Ix\leq\bm{0}$.
Let $D$ be as in Theorem 1 depending on $\Delta$ and $n$ (and thus $m$) only. Now, fix any $r\in\{0,\ldots,D-1\}^{m}$ and let $B_{r},C_{r},f_{r}$ be the vectors so that
$P(b)_{I}=\{x\in\mathbb{R}^{n}:B_{r}x\leq f_{r}+C_{r}b\}$
for all $b\in\mathbb{Z}^{m}$ with $b\equiv r\pmod{D}$. Let $Q_{r}=\{b\in\mathbb{R}^{m}\mid\exists x\in\mathbb{R}^{n}:Bx\leq f+Cb\}$. Then $Q_{r}$ is the linear projection of a polyhedron and hence it is again a polyhedron.
Now, consider a right hand side $b\in\mathbb{Z}^{m}$ with $b\equiv r\pmod{D}$. Then
$b\in\operatorname{intcone}(W)$$\Leftrightarrow$ $P(b)\cap\mathbb{Z}^{n}\neq\emptyset$
$\Leftrightarrow$ $P(b)_{I}\neq\emptyset$
$\Leftrightarrow$ $b\in Q_{r}.$
4.2 Optimizing over 2-stage stochastic IPs
A 2-stage stochastic integer program is of the form
$\max c^{T}x+\sum_{i=1}^{n}d_{i}^{T}y_{i}$ $(\textsc{2SSIP})$
$U_{i}x+V_{i}y_{i}$ $=$ $b_{i}\quad\forall i=1,\ldots,n$
$x,y_{1},\ldots,y_{n}$$\in$ $\mathbb{Z}_{\geq 0}^{k},$
[…]
Theorem 7.
The problem $(\textsc{2SSIP})$ can be solved in time $g(k,\Delta)$ times a fixed polynomial in the encoding length where $\Delta=\max\{\|V_{i}\|_{\infty}:i=1,\ldots,n\}$.
Proof.
We use the following algorithm:
  1. (1)
    Let $P_{i}(b^{\prime})=\{y_{i}\in\mathbb{R}^{k}\mid V_{i}y_{i}=b^{\prime},y_{i}\geq\bm{0}\}$ for $i=1,\ldots,n$.
  2. (2)
    Let $D$ be the value as in Theorem 1 that works for parameters $k$ and $\Delta$.
  3. (3)
    Guess $r\in\{0,\ldots,D-1\}^{k}$ so that $x^{**}\equiv r\pmod{D}$ where $(x^{**},y_{1}^{**},\ldots,y_{n}^{**})$ is an optimum integral solution to $(\textsc{2SSIP})$.
  4. (4)
    For all $i=1,\ldots,n$, let $r_{i}\in\{0,\ldots,D-1\}^{k}$ with $r_{i}=b_{i}-U_{i}r\pmod{D}$.
  5. (5)
    For each $i\in[n]$, apply Theorem 1 to write $P_{i}(b^{\prime})_{I}=\{y_{i}\in\mathbb{R}^{k}\mid B_{i}y_{i}\leq f_{i}+C_{i}b^{\prime}\}$ for all $b^{\prime}$ with $b^{\prime}\equiv r_{i}\pmod{D}$.
  6. (6)
    Compute an optimum extreme point solution $(x^{*},y_{1}^{*},\ldots,y_{n}^{*})$ to the mixed integer linear program
    $\max c^{T}x+\sum_{i=1}^{n}d_{i}^{T}y_{i}$
    $B_{i}y_{i}$$\leq$ $f_{i}+C_{i}(b_{i}-U_{i}x)$
    $x$$\equiv$ $r\pmod{D}$
    $x$$\in$ $\mathbb{Z}_{\geq 0}^{k}$
  7. (7)
    Return $(x^{*},y_{1}^{*},\ldots,y_{n}^{*})$
We know that $b_{i}-U_{i}x^{**}\equiv r_{i}$ and so $(x^{**},y_{1}^{**},\ldots,y_{n}^{**})$ is feasible for the MIP in (6). The MIP in $(6)$ has only $k$ many integral variables (and $nk$ many fractional ones) and hence it can be solved within the claimed running time. On the other hand, if $(x^{*},y_{1}^{*},\ldots,y_{n}^{*})$ is an optimum extreme point to (6), then $x^{*}\in\mathbb{Z}_{\geq 0}^{k}$ and each $y^{**}_{i}$ is an extreme point to $P_{i}(b_{i}-U_{i}x^{*})_{I}$. Each such extreme point must be integral. That shows the claim. ∎
4.3 An almost solution to the 4-block problem
The 4-block problem is the IP of the form
$\max c^{T}x+\sum_{i=1}^{n}d_{i}^{T}y_{i}$ $({\textsc{4BlockIP}})$
$Wx+X_{1}y_{1}+\ldots+X_{n}y_{n}$ $=$ $a\quad(*)$
$U_{i}x+V_{i}y_{i}$ $=$ $b_{i}\quad\forall i=1,\ldots,n$
$x,y_{1},\ldots,y_{n}$$\in$ $\mathbb{Z}_{\geq 0}^{k}$
where $U_{i},V_{i},W,X_{i}\in\mathbb{Z}^{k\times k}$ and $a,b_{i}\in\mathbb{Z}^{k}$. This is a strict generalization of $({\textsc{2SSIP}})$. Note that after deleting $k$ many variables and $k$ many constraints, the problem decomposes into disjoint IPs with $k$ variables each. Indeed, $(\textsc{4BlockIP})$ can be solved in time $n^{f(k,\Delta)}$ times a fixed polynomial in the encoding length where $\Delta=\max\{\|U_{i}\|_{\infty},\|V_{i}\|_{\infty},\|W\|_{\infty},\|X_{i}\|_{\infty}:i\in[n]\}$, see the work of Hemmecke, Köppe and Weismantel [13]. It is a popular open problem in the theoretical IP community whether there is an FPT-type algorithm as well.
Conjecture 1 (4-block conjecture).
$(\textsc{4BlockIP})$ can be solved in time $f(k,\Delta)$ times a fixed polynomial in the encoding length where $\Delta$ is the largest absolute value appearing in $U_{i},V_{i},W,X_{i}$.
We prove that one can at least find an almost feasible solution that violates the $O(1)$ many joint constraints by an additive $O(1)$ (assuming $k$ and the maximum entries are bounded by constants):
Theorem 8.
Suppose that $(\textsc{4BlockIP})$ is feasible with value $OPT$ and let $\Delta=\max\{\|V_{i}\|_{\infty}:i\in[n]\}$. Then in time $f(k,\Delta)$ times a polynomial in the input length one can find a vector $(x^{*},y_{1}^{*},\ldots,y_{n}^{*})$ with $c^{T}x^{*}+\sum_{i=1}^{n}d_{i}^{T}y_{i}^{*}\geq OPT$ that satisfies all constraints except ()(*)( ∗ ). Moreover, the error for constraint ()(*)( ∗ ) is bounded by
$\big{\|}(Wx^{*}+X_{1}y_{1}^{*}+\ldots+X_{n}y_{n}^{*})-a\big{\|}_{\infty}\leq g(k,\max_{i=1,\ldots,n}\|X_{i}\|_{\infty})$
Proof.
We use a similar argument to Theorem 7. Again set $P_{i}(b^{\prime})=\{y_{i}\in\mathbb{R}^{k}\mid V_{i}y_{i}=b^{\prime},y_{i}\geq\bm{0}\}$. Let $(x^{**},y_{1}^{**},\ldots,y_{n}^{**})$ be a solution to $({\textsc{4BlockIP}})$ attaining the value of $OPT$. Again we can guess a vector $r$ so that $x^{**}\equiv r\pmod{D}$ where $D$ is a parameter so that by Theorem 1 one has
$P_{i}(b^{\prime})_{I}=\{y_{i}\in\mathbb{R}^{k}\mid B_{i}y_{i}\leq f_{i}+C_{i}b^{\prime}\}$
for all $b^{\prime}$ with $b^{\prime}\equiv r_{i}\pmod{D}$ where $r_{i}=b_{i}-U_{i}r\pmod{D}$. Consider the mixed integer linear program
$\max c^{T}x+\sum_{i=1}^{n}d_{i}^{T}y_{i}$ $({\textsc{4BlockMIP}})$
$Wx+X_{1}y_{1}+\ldots+X_{n}y_{n}$ $=$ $a\quad(**)$
$y_{i}$$\in$ $P_{i}(b_{i}-U_{i}x)_{I}\quad\forall i=1,\ldots,n$
$x$$\equiv$ $r\pmod{D}$
$x$$\in$ $\mathbb{Z}_{\geq 0}^{k}$
We compute an optimum extreme point solution $(x^{*},\tilde{y}_{1},\ldots,\tilde{y}_{n})$ to $({\textsc{4BlockMIP}})$. We note that $x^{*}$ will be integral, but as all vertices of $P_{i}(\cdot)_{I}$ are integral and ()(**)( ∗ ∗ ) are only $k$ non-trivial constraints, the set $J=\{i\in[n]\mid\tilde{y}_{i}\notin\mathbb{Z}^{k}\}$ has size $|J|\leq k$. Then by standard proximity bounds, for each $i\in J$, there is a $y_{i}^{*}\in P_{i}(b_{i}-U_{i}x^{*})_{I}\cap\mathbb{Z}^{k}$ with $\|y_{i}^{*}-\tilde{y}_{i}\|_{\infty}\leq(\Delta k)^{O(k)}$ and $d_{i}^{T}y_{i}^{*}\geq d_{i}^{T}\tilde{y}_{i}$. We set $y_{i}^{*}=\tilde{y}_{i}$ for all $i\notin J$ and $(x^{*},y_{1}^{*},\ldots,y_{n}^{*})$ satisfies the claim. ∎